OpenStudy (anonymous):

Which set of side lengths will form a right triangle? 6, 8, 11 10, 24, 25 7, 24, 26 8, 15, 17

Nnesha (nnesha):

use Pythagorean theorem equation \[\huge\rm a^2+b^2=c^2\] c= longest side if you get equal sides then those 3 numbers form a right triangle

OpenStudy (anonymous):

c

Nnesha (nnesha):

how ?

OpenStudy (anonymous):

how u get that?